30y to the fourth power divided by -5y to the second power

Found 2 solutions by dabanfield, Fombitz:
Answer by dabanfield(803)

30y to the fourth power divided by -5y to the second power

We have 30y^4/-5y so:
30y^4/-5y = (30/-5)*(y^4)*(y^(-1)) = -6*y^(4-1) = -6*y^3

Answer by Fombitz(32388)